Danny invested some money at 3% simple interest in some money at 7% simple interest. The amount invested at the higher rate was $5000 more than the amount invested at the lower rate. If the total interest on the investments for one year was $834 and how much did he invest at each rate?
Answer by lwsshak3(11628)

Danny invested some money at 3% simple interest in some money at 7% simple interest. The amount invested at the higher rate was $5000 more than the amount invested at the lower rate. If the total interest on the investments for one year was $834 and how much did he invest at each rate?
***
let x=amt invested at 3%
x+5000=amt invested at 7%
3%x+7%(x+5000)=834
.03x+.07x+350=834
.10x=484
x=4840
x+5000=9840
amt invested at 3%=$4840
amt invested at 7%=$9840